# Soft Theorems in Chern-Simons Matter Theories

### v3

@@ 7 Conclusion
(1) the subleading theorem gets modified when Chern-Simons terms are added and therefore, it is no longer universal even though the theory is still gauge-invariant. Also, such terms cannot be obtained from Sen’s covariantisation procedure and need to be added separately.
(2) we can in principle take arbitrary values of $\left\{c_{4},c_{5}\right\}$ which would leave $\mathcal{S}_{CS}$ completely non gauge-invariant and still the amplitudes will obey the "universal" leading soft theorem. Therefore, we have an example of a non gauge-invariant action that still satisfies the usual universal leading soft theorem at the tree level. This means that even though gauge invariance implies the universal form of leading soft theorems, the opposite is not true. There are examples of non-gauge invariant theories which still obey the same leading soft theorem. I have provided one such example.
@@ Acknowledgments
I would like to thank my PhD advisor, R. Loganayagam, for helping me with the initial idea and in setting up the problem as well as for countless valuable insights. I would also like to thank Biswajit Sahoo and Ashoke Sen for their invaluable feedback on my draft. Finally, I would like to thank my colleagues Godwin Martin, Kaustubh Singhi and Shridhar Vinayak for contributing to various discussions throughout this work. I also acknowledge the support of the Department of Atomic Energy, Government of India, under project no. RTI4001.

### v4

@@ 8 Conclusion
Firstly, the subleading term gets modified when Chern-Simons terms are added and therefore, it is no longer universal even though the amplitude is still gauge-invariant under small gauge transformations. Also, such terms cannot be obtained from Sen’s covariantisation procedure and need to be added separately.
Secondly, we can in principle take arbitrary values of $\left\{c_{4},c_{5}\right\}$ which would leave the deformation $S_{CS}$ completely non-gauge-invariant. Even under such non-gauge-invariant deformations, the amplitudes still obey the "universal" leading soft theorem, at least at the tree-level. Therefore, we have an example of a non-gauge-invariant action that still satisfies the usual universal leading soft theorem at the tree-level. This means that gauge invariance of amplitudes is a sufficient but not necessary condition for the universal form of the tree-level leading soft theorem.
The analysis in this paper is limited to tree-level amplitudes. Calculations at higher loop orders may modify the second statement since the loss of gauge invariance should imply fewer constraints on the amplitudes. Therefore, one would expect that to show up in the IR behaviour of the amplitudes at higher loop orders.
@@ Acknowledgments
I would like to thank, R. Loganayagam, for helping me with the initial idea and in setting up the problem as well as for countless valuable insights. I would also like to thank Biswajit Sahoo and Ashoke Sen for their invaluable feedback on the draft. I would also like to thank the referee for their insightful comments on the manuscript. Finally, I would like to thank my colleagues Godwin Martin, Kaustubh Singhi and Shridhar Vinayak for contributing to various discussions throughout this work. I also acknowledge the support of the Department of Atomic Energy, Government of India, project no. RTI4001.